OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? A) 2 B) 3/4 C)-4/3 D) -2

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we have to do this computation: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{3 - \left( { - 7} \right)}}{{ - 1 - 4}} = ...?\]

OpenStudy (michele_laino):

being \((x_1,y_1)=(4,-7)\), and \((x_2,y_2)=(-1,3)\)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

the answer is d(-2)